OpenStudy (anonymous):

3+logsubscript3 (x+3)=6

OpenStudy (zzr0ck3r):

\(3+\log_3(x+3)=6\implies \log_3(x+3)=3\implies3^3=x+3\implies x=24\)

OpenStudy (zzr0ck3r):

all we use is the property that \(\large\log_a(b)=c \iff a^c=b\)

OpenStudy (anonymous):

Thank you, I haven't done logs in a few years. I need to brush up on it.